OpenStudy (anonymous):

The perimeter of the rectangle is fifty-four feet. What is x? A rectangle is shown. The length is labeled fifteen feet and width is labeled x.

OpenStudy (texaschic101):

P = 2L + 2W.....we will go ahead and label x for W P = 2L + 2x P = 54 L = 15 now we sub 54 = 2(15) + 2x 54 = 30 + 2x 54 - 30 = 2x 24 = 2x 12 = x check.. P = 2L + 2x 54 = 2(15) + 2(12) 54 = 30 + 24 54 = 54 (correct) the width (x) = 12 cm

OpenStudy (anonymous):

thank you are a big help

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

have any other questions